# Anticipating Gaming to Incentivize Improvement: Guiding Agents in (Fair) Strategic Classification

### v1

@@ A.2 Proofs for Section 5
It remains to find the specific forms of the first derivatives $\bm{\Phi}_{s,(i)}^{{}^{\prime}y}$ and $\bm{\Psi}_{s,(i)}^{{}^{\prime}y}$ for each equilibrium type (i).
We show this for agents’ best-response Type 1, finding the expressions in the first row in Table 3. A similar procedure can be used to find the unfair strategic decision threshold for agents’ best-response Type 2 and 3.
@@
From the definition of the indifference points in Definition 1, we have $\mathbb{T}^{y}_{I,s}({\theta_{s}}-\mathbf{o}^{y}_{I,s})=\mathbb{T}^{y}_{I,s}((\mathbb{T}^{y}_{I,s})^{-1}(1-C_{I,s}))=1-C_{I,s}$. We also know, from Proposition 2, that the conditions for this equilibrium to occur are that $\mathbf{f}^{y}_{s}\leq\mathbf{o}^{y}_{I,s}\leq\mathbf{o}^{y}_{M,s}\leq\mathbf{r}^{y}_{s}$, and therefore, $11-\mathbb{T}^{y}_{I,s}({\theta_{s}}-\mathbf{r}^{y}_{s})=$ 1 (i.e., admission under $I$ is certain at $\mathbf{r}^{y}_{s}$). Therefore:
@@
$\frac{\partial\mathbf{f}^{y}_{s}}{\partial\theta_{s}}=1-\frac{\partial h(g(\theta_{s},\mathbf{f}^{y}_{s}))}{\partial\theta_{s}}.\text{ Next, the derivative of $h(g(\theta_{s},\mathbf{f}^{y}_{s}))$ with respect to $\theta_{s}$ is given by:}$:
@@
$\frac{\partial h(g(\theta_{s},\mathbf{f}^{y}_{s}))}{\partial\theta_{s}}=\frac{\tau^{y}_{I,s}(\theta_{s}-\mathbf{f}^{y}_{s})}{\tau^{y}_{M,s}(\theta_{s}-\mathbf{f}^{y}_{s})}\cdot\big{(}1-\frac{\partial\mathbf{f}^{y}_{s}}{\partial\theta_{s}}\big{)}.\text{ Substituting this back into the expression for $\frac{\partial\mathbf{f}^{y}_{s}}{\partial\theta_{s}}$},\text{ we get:}$:
@@
By setting the first derivative to zero and re-arrange terms we have $\theta_{s}^{\mathcal{C}}$ satisfy: $\sum_{s}n_{s}\frac{u_{+}\alpha_{s}G^{1}_{s}(\theta_{s}^{\mathcal{C}})-u_{-}(1-\alpha_{s})G^{0}_{s}(\theta_{s}^{\mathcal{C}})}{\frac{\partial{\mathcal{C}^{f}_{s}(\theta_{s}^{\mathcal{C}})}}{\partial{\theta_{s}^{\mathcal{C}}}}}=0$.
A similar approach can prove Lemma 7 using the strategic utility from Eq. 5. We omit it due to space constraints.
A.3 Proof of Proposition 8
For this part of the proof, we analyze the selected optimal unfair thresholds ($\theta^{U}_{s},\hat{\theta}^{U}_{s}$) over different groups’ qualification status ($\alpha_{s}$) and agents’ best response types (Type 1 and 3). Let $\alpha_{s}=0$ such that all agents are unqualified, the unfair non-strategic firm’s utility in equation 2 can be reduced to $U(\theta_{s})=-u_{-}(1-\mathbb{G}^{0}_{s}(\theta_{s}))$, hence to achieve the maximum utility the optimal threshold is $\theta_{s}=\overline{x}^{0}_{s}$ which means rejecting all the unqualified agents. Moreover, Let $\alpha_{s}=1$ such that all agents are qualified, the unfair non-strategic firm’s utility in equation 2 can be reduced to $U(\theta_{s})=u_{+}(1-\mathbb{G}^{1}_{s}(\theta_{s}))$, hence the optimal threshold is $\theta_{s}=\underline{x}^{1}_{s}$ which means accepting majority all qualified agents. Note as the group qualification rate ($\alpha_{s}$) decreases, the unfair non-strategic optimal threshold increases to reject the more unqualified agents. To summarize, the unfair non-strategic optimal policy satisfies the following: $[\underline{x}^{1}_{s},\overline{x}^{0}_{s}]$. Under Assumption 4, $\mathbf{o}^{y}_{M,s}$ and $\mathbf{r}^{y}_{s}$ are equal across agents, so we denote them as $\mathbf{o}_{M,s}$ and $\mathbf{r}_{s}$, omitting the $y$ label in this proof. To establish a unique optimal strategic utility under Type 3 agent responses, we analyze the derivative from Table 3:
$\frac{\partial{\hat{U}(\hat{\theta}^{U}_{s})}}{\partial{\hat{\theta}^{U}_{s}}}=u_{+}\alpha_{s}\left(-G^{1}_{s}(\mathbf{o}_{M,s})C_{M,s}-(G^{1}_{s}*\tau_{M,s})(\hat{\theta}^{U}_{s})\right)+u_{-}(1-\alpha_{s})\left(G^{0}_{s}(\mathbf{o}_{M,s})C_{M,s}+(G^{0}_{s}*\tau_{M,s})(\hat{\theta}^{U}_{s})\right).$
@@ A.3 Proof of Proposition 8
We can re-write $\bm{\Psi}^{y}_{s,(3)}(\theta_{s})=\int_{\mathbf{o}_{M,s}}^{\theta_{s}}h^{y}_{M,s}(z;\theta)\mathrm{d}z$, and because the integration limits depends on $\theta_{s}$, the following holds: $\text{If }\mathbf{o}_{M,s}<\theta_{s}\leq z^{*},\int_{\mathbf{o}_{M,s}}^{\theta_{s}}h^{y}_{M,s}(z;\theta_{s})\mathrm{d}z\text{ increases as $\theta_{s}$ approaches the $z^{*}$},$ and $\text{If }z^{*}\leq\mathbf{o}_{M,s}<\theta_{s},\int_{\mathbf{o}_{M,s}}^{\theta_{s}}h^{y}_{M,s}(z;\theta_{s})\mathrm{d}z\text{ decreases as $\mathbf{o}_{M,s}$ moves away from the $z^{*}$}.$ Therefore, $\bm{\Psi}^{y}_{s,(3)}(\theta_{s})$ is also single-peaked. Similarly, $\bm{\Psi}^{y}_{s,(i)}(\theta_{s})\text{ for }i\in\{1,2,3\}$, and $\bm{\Phi}^{y}_{s,(i)}(\theta_{s})\text{ for }i\in\{1,2\}$ are also single peaked.
@@
Given the analysis of optimal thresholds, we first order the policies as in Proposition 8 and then rank their impacts as per parts (i), (ii), and (iii) of the same proposition. The maximum value of the optimal unfair strategic threshold, $\hat{\theta}^{U}_{s}$, under Type 1 agent responses is $\hat{\theta}^{U}_{s}\leq\overline{x}^{0}_{s}+(\mathbb{T}^{0}_{M,s})^{-1}(C_{I,s}-C_{M,s})<\mu^{0}_{s}+\underline{b}^{0}_{I,s}<\mu^{1}_{s}$. Thus, $\hat{\theta}^{U}_{s}$ is always less than $\arg\max\bm{\Phi}^{y}_{s,(1)}(\theta_{s})$ and $\arg\max\bm{\Psi}^{1}_{s,(1)}(\theta_{s})$, placing it in the first half of both $\bm{\Phi}^{y}_{s,(1)}(\theta_{s})$ and $\bm{\Psi}^{1}_{s,(1)}(\theta_{s})$ The minimum $\hat{\theta}^{U}_{s}>\mu^{0}_{s}+\underline{b}_{M,s}>\arg\max\bm{\Psi}^{0}_{s,(1)}(\theta_{s})$ (second half of $\bm{\Psi}^{0}_{s,(1)}(\theta_{s}))$.
For the optimal non-strategic threshold, $\theta^{U}_{s}<\overline{x}^{0}_{s}<\mu^{0}_{s}+\underline{b}^{0}_{I,s}<\mu^{1}_{s}$, it is always less than $\arg\max\bm{\Phi}^{y}_{s,(1)}(\theta_{s})$ and $\arg\max\bm{\Psi}^{1}_{s,(1)}(\theta_{s})$, residing in the first half of $\bm{\Phi}^{y}_{s,(1)}(\theta_{s})$ and $\bm{\Psi}^{1}_{s,(1)}(\theta_{s})$. Its minimum, $\theta^{U}_{s}>\underline{x}^{1}_{s}>\mu^{0}_{s}+\underline{b}_{M,s}$, exceeds $\arg\max\bm{\Psi}^{0}_{s,(1)}(\theta_{s})$, placing it in the second half of $\bm{\Psi}^{0}_{s,(1)}(\theta_{s})$. Since the optimal strategic threshold always exceeds the non-strategic one, and both lie in the first half of $\bm{\Phi}^{y}_{s,(1)}(\theta_{s})$ and $\bm{\Psi}^{1}_{s,(1)}(\theta_{s})$, their impacts can be ordered as $\bm{\Phi}^{y}_{s,(1)}(\theta^{U}_{s})<\bm{\Phi}^{y}_{s,(1)}(\hat{\theta}^{U}_{s})$ and $\bm{\Psi}^{1}_{s,(1)}(\theta^{U}_{s})<\bm{\Psi}^{1}_{s,(1)}(\hat{\theta}^{U}_{s})$. Additionally, as both reside in the second half of $\bm{\Psi}^{0}_{s,(1)}(\theta_{s})$, their impacts order as $\bm{\Psi}^{0}_{s,(1)}(\hat{\theta}^{U}_{s})<\bm{\Psi}^{0}_{s,(1)}(\theta^{U}_{s})$.
@@
A.4 Impacts of relaxing the assumptions of Proposition 8

### v2

@@ A.2.2 Proof of Lemma 5
It remains to find the specific forms of the first derivatives $\bm{\Phi}_{s,(i)}^{{}^{\prime}y}$ and $\bm{\Psi}_{s,(i)}^{{}^{\prime}y}$ for each equilibrium type (i). We show this for agents’ best-response Type 1, finding the expressions in the first row in Table 2. A similar procedure can be used to find the unfair strategic decision threshold for agents’ best-response Type 2 and 3.
@@
From the definition of the indifference points in Definition 1, we have $\mathbb{T}^{y}_{I,s}({\theta_{s}}-\mathbf{o}^{y}_{I,s})=\mathbb{T}^{y}_{I,s}((\mathbb{T}^{y}_{I,s})^{-1}(1-C_{I,s}))=1-C_{I,s}$. We also know, from Proposition 2, that the conditions for this equilibrium to occur are that $\mathbf{f}^{y}_{s}\leq\mathbf{o}^{y}_{I,s}\leq\mathbf{o}^{y}_{M,s}\leq\mathbf{r}^{y}_{s}$, and therefore, $1-\mathbb{T}^{y}_{I,s}({\theta_{s}}-\mathbf{r}^{y}_{s})=1$ (i.e., admission under $I$ is certain at $\mathbf{r}^{y}_{s}$). Therefore:
@@
[…]
@@
[…]
@@ A.2.3 Proof of Lemmas 6-7
[…]
A similar approach can prove Lemma 7 using the strategic utility from (5).
A.3 Proofs for Section 6
A.3.1 Proof of Proposition 8
For this part of the proof, we analyze the selected optimal unfair thresholds ($\theta^{U}_{s},\hat{\theta}^{U}_{s}$) over different groups’ qualification status ($\alpha_{s}$) and agents’ best response types (Type 1 and 3). Let $\alpha_{s}=0$ such that all agents are unqualified, the unfair non-strategic firm’s utility in equation 2 can be reduced to $U(\theta_{s})=-u_{-}(1-\mathbb{G}^{0}_{s}(\theta_{s}))$, hence to achieve the maximum utility the optimal threshold is $\theta_{s}=\overline{x}^{0}_{s}$ which means rejecting all the unqualified agents. Moreover, Let $\alpha_{s}=1$ such that all agents are qualified, the unfair non-strategic firm’s utility in equation 2 can be reduced to $U(\theta_{s})=u_{+}(1-\mathbb{G}^{1}_{s}(\theta_{s}))$, hence the optimal threshold is $\theta_{s}=\underline{x}^{1}_{s}$ which means accepting majority all qualified agents. Note as the group qualification rate ($\alpha_{s}$) decreases, the unfair non-strategic optimal threshold increases to reject the more unqualified agents. To summarize, the unfair non-strategic optimal policy satisfies the following: $\underline{x}^{1}_{s},\overline{x}^{0}_{s}]$. Under Assumption 4, $\mathbf{o}^{y}_{M,s}$ and $\mathbf{r}^{y}_{s}$ are equal across agents, so we denote them as $\mathbf{o}_{M,s}$ and $\mathbf{r}_{s}$, omitting the $y$ label in this proof. To establish a unique optimal strategic utility under Type 3 agent responses, we analyze the derivative from Table 2:
$\frac{\partial{\hat{U}(\hat{\theta}^{U}_{s})}}{\partial{\hat{\theta}^{U}_{s}}}=$ $u_{+}\alpha_{s}\left(-G^{1}_{s}(\mathbf{o}_{M,s})C_{M,s}-(G^{1}_{s}*\tau_{M,s})(\hat{\theta}^{U}_{s})\right)$
$+u_{-}(1-\alpha_{s})\left(G^{0}_{s}(\mathbf{o}_{M,s})C_{M,s}+(G^{0}_{s}*\tau_{M,s})(\hat{\theta}^{U}_{s})\right).$
@@ A.3.1 Proof of Proposition 8
We can re-write $\bm{\Psi}^{y}_{s,(3)}(\theta_{s})=\int_{\mathbf{o}_{M,s}}^{\theta_{s}}h^{y}_{M,s}(z;\theta)\mathrm{d}z$, and because the integration limits depends on $\theta_{s}$, the following holds: If $\mathbf{o}_{M,s}<\theta_{s}\leq z^{*},\int_{\mathbf{o}_{M,s}}^{\theta_{s}}h^{y}_{M,s}(z;\theta_{s})\mathrm{d}z$ increases as $\theta_{s}$ approaches the $z^{*}$, and If $z^{*}\leq\mathbf{o}_{M,s}<\theta_{s},\int_{\mathbf{o}_{M,s}}^{\theta_{s}}h^{y}_{M,s}(z;\theta_{s})\mathrm{d}z$ decreases as $\mathbf{o}_{M,s}$ moves away from the $z^{*}$. Therefore, $\bm{\Psi}^{y}_{s,(3)}(\theta_{s})$ is also single-peaked. Similarly, $\bm{\Psi}^{y}_{s,(i)}(\theta_{s})$ for $i\in\{1,2,3\}$, and $\bm{\Phi}^{y}_{s,(i)}(\theta_{s})$ for $i\in\{1,2\}$ are also single peaked.
@@
Given the analysis of optimal thresholds, we first order the policies as in Proposition 8 and then rank their impacts as per parts (i), (ii), and (iii) of the same proposition. The maximum value of the optimal unfair strategic threshold, $\hat{\theta}^{U}_{s}$, under Type 1 agent responses is $\hat{\theta}^{U}_{s}\leq\overline{x}^{0}_{s}+(\mathbb{T}^{0}_{M,s})^{-1}(C_{I,s}-C_{M,s})<\mu^{0}_{s}+\underline{b}^{0}_{I,s}<\mu^{1}_{s}$. Thus, $\hat{\theta}^{U}_{s}$ is always less than $\arg\max\bm{\Phi}^{y}_{s,(1)}(\theta_{s})$ and $\arg\max\bm{\Psi}^{1}_{s,(1)}(\theta_{s})$, placing it in the first half of both $\bm{\Phi}^{y}_{s,(1)}(\theta_{s})$ and $\bm{\Psi}^{1}_{s,(1)}(\theta_{s})$. The minimum $\hat{\theta}^{U}_{s}>\mu^{0}_{s}+\underline{b}_{M,s}>\arg\max\bm{\Psi}^{0}_{s,(1)}(\theta_{s})$ (second half of $\bm{\Psi}^{0}_{s,(1)}(\theta_{s}))$.
[…]
@@
A.3.2 Impacts of relaxing the assumptions of Proposition 8